A certain natural decay series stars with 92238U and ends with 82204Pb. Each step involves the loss of either an α-particle or a β-particle. What is the number of α-particles released?

What is the number of β-particles released?



1
Expert's answer
2021-09-29T05:03:07-0400

α-particles = 238-204 = 34

β-particle = 92-82 = 10